OpenStudy (theedwardsfamily):

A lil' help please? :P The altitude of a plane as it begins to descend is shown by the linear equation a = 22,000 – 800m, where a is the plane’s altitude, in feet, and m is minutes. What is the plane’s altitude when m = 12? _____ ft

OpenStudy (theedwardsfamily):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

substitute 12 for m and solve

OpenStudy (alexandervonhumboldt2):

a=22000-800-12=?

OpenStudy (alexandervonhumboldt2):

woops

OpenStudy (alexandervonhumboldt2):

i mean a=22000-800*12

OpenStudy (theedwardsfamily):

lol

OpenStudy (theedwardsfamily):

i got 254400

OpenStudy (alexandervonhumboldt2):

how?

OpenStudy (alexandervonhumboldt2):

22000-9600=?

OpenStudy (theedwardsfamily):

ohhh lol one sec

OpenStudy (theedwardsfamily):

I subtracted first xD

OpenStudy (theedwardsfamily):

i got 12400

OpenStudy (alexandervonhumboldt2):

yeah now it is correct

OpenStudy (kfireangel):

\[A=22000-(800*12)=\]

OpenStudy (theedwardsfamily):

Yas! thnx :P